The perimeter of a yield sign is 22 feet and has the shape of an isosceles triangle. If the shortest side is two feet less than the other two sides, find the length of the shortest side.
I need to set up an algebraic expression to solve problem

Answer by rothauserc(4718)

let x be the shortest side and y the length of one of the other equal sides,then
y = x + 2
Perimeter(P) = x + 2y
22 = x + 2*(x+2)
22 = x + 2x + 4
3x = 18
x = 6
the shortest side is 6 feet and the other two sides are 8 feet each